# Inhomogeneous Kinetic Equation for Mixed Neutrinos: Tracing the Missing Energy

### v1

@@ Abstract
Flavor-dependent neutrino transport is described by a well-known kinetic equation for occupation-number matrices in flavor space. However, as an overlooked theoretical problem, we show that in the inhomogeneous case, neutrino-neutrino refractive energy is not conserved. We derive the missing gradient terms in the fast flavor limit (vanishing neutrino masses), and prove that the missing refractive energy is traded with the huge reservoir of neutrino kinetic energy through gradients of the weak interaction potential. Even small changes of the kinetic energy accommodate the refractive energy gained or lost. Flavor evolution alone is negligibly affected by the new terms.
Introduction.—Flavor evolution in a dense neutrino environment is often described by a kinetic equation for the flavor density matrices $\rho({\bf p},{\bf x},t)$, where the diagonal entries are the usual occupation numbers Dolgov:1980cq ; Rudsky ; Sigl:1993ctk ; Sirera:1998ia ; Yamada:2000za ; Vlasenko:2013fja ; Volpe:2013uxl ; Serreau:2014cfa ; Kartavtsev:2015eva ; Fiorillo:2024wej . In the homogeneous case, they are defined as the expectation values of the number-operator matrices ${\cal D}_{\alpha\beta}=a^{\dagger}_{\beta,{\bf p}}a_{\alpha,{\bf p}}$ with the flavor indices $\alpha$ and $\beta$ and $a^{\dagger}_{\alpha,{\bf p}}$ the creation operator for a left-handed neutrino of flavor $\alpha$ with momentum ${\bf p}$. In the limit of vanishing neutrino masses and flavor mixing, nontrivial flavor evolution can still arise through classical instabilities that would engender strong flavor correlations Sawyer:2015dsa ; Chakraborty:2016lct ; Izaguirre:2016gsx ; Tamborra:2020cul ; Richers:2022zug . Dubbed fast flavor evolution, this subject has been intensely studied in view of practical consequences for neutrino flavor transport in stellar core collapse and binary neutron star mergers Xiong:2022vsy ; Fernandez:2022yyv ; Ehring:2023abs ; Ehring:2023lcd ; Nagakura:2023mhr ; Nagakura:2023xhc ; Zaizen:2023wht ; Grohs:2023pgq ; Abbar:2023zkm ; Akaho:2023brj ; Froustey:2023skf ; Nagakura:2023wbf .
Instead, we turn to an overlooked theoretical problem of the usual equations of motion (EOMs): their lack of energy conservation. In their often-used form, they are
@@
where ${\sf H}_{{\bf p},{\bf r},t}=\sqrt{2}G_{\rm F}\sum_{{\bf p}^{\prime}}\rho_{{\bf p}^{\prime},{\bf r},t}\,(1-\cos\theta_{{\bf p},{\bf p}^{\prime}})$ is a matrix that drives flavor evolution by neutrino-neutrino refraction. The so-called advection term proportional to ${\bf v}={\bf p}/|{\bf p}|$ quantifies a drift in coordinate space caused by inhomogeneities. Among other consequences, this term allows for ${\bf p}$-dependent instabilities and even an initially homogeneous system can develop flavor disturbances that grow exponentially and drift to ever smaller scales Duan:2014gfa ; Mirizzi:2015fva ; Bhattacharyya:2020jpj ; Johns:2020qsk ; Tamborra:2020cul ; Richers:2021nbx ; Bhattacharyya:2022eed .
@@
Interaction energy.—We first need to define the $\nu$$\nu$ interaction energy. The starting point is the many-body Hamiltonian $\mathcal{H}=\mathcal{H}_{0}+\mathcal{U}$, where $\mathcal{H}_{0}=\sum_{\alpha,{\bf p}}\epsilon_{\bf p}a^{\dagger}_{\alpha,{\bf p}}a_{\alpha,{\bf p}}$ is the kinetic energy operator with $\epsilon_{\bf p}=|{\bf p}|$ in the massless limit. Moreover, the interaction energy is
@@
The energy of the system is the expectation value of $\mathcal{H}$ over the quantum state. In particular, the kinetic energy is $K=\langle\mathcal{H}_{0}\rangle=\int d^{3}{\bf r}\sum_{\bf p}\epsilon_{\bf p}\mathrm{Tr}(\rho_{\bf p})$. For the interaction energy, we need to determine the average value of a string of four creation and annihilation operators. We do so in the mean field, or Hartree-Fock, approximation, where the quantum state is assumed to be the product of single-particle states, so that the expectation value factorizes as $\langle a^{\dagger}_{1}a_{2}a^{\dagger}_{3}a_{4}\rangle=\langle a^{\dagger}_{1}a_{2}\rangle\langle a^{\dagger}_{3}a_{4}\rangle+(\delta_{2,3}-\langle a^{\dagger}_{3}a_{2}\rangle)\langle a^{\dagger}_{1}a_{4}\rangle$. In this way we define a potential energy of interaction $U=\langle\mathcal{U}\rangle$. For a homogeneous system, it is
@@
where we neglect a single-particle term constituting an irrelevant (infinite) renormalization of the neutrino chemical potential.
@@
Different from a neutrino gas subject to collisions with matter (see, e.g., Refs. Johns:2021qby ; Nagakura:2022qko ; Xiong:2022zqz ; Fiorillo:2024 ; Liu:2023pjw ; Lin:2022dek ; Johns:2022yqy ; Shalgar:2022rjj ; Padilla-Gay:2022wck ; Fiorillo:2023ajs ; Johns:2023xae ), where energy is not conserved anyway, Eqs. (Inhomogeneous Kinetic Equation for Mixed Neutrinos: Tracing the Missing Energy) describe a closed system that admits a conserved energy. The kinetic part $K=\int d^{3}{\bf r}\sum_{\bf p}\epsilon_{\bf p}P^{0}_{\bf p}$ is seen to be conserved from Eq. (6a). Also the nonoscillatory part of the interaction energy is separately conserved. $U_{\mathrm{osc}}=(\sqrt{2}G_{\rm F}/4)\int d^{3}{\bf r}\,({\vec{P}}_{0}^{2}-{\vec{P}}_{1}^{2})$, which follows from Eq. (5), must then be conserved as well. In the homogeneous case, this is indeed the case; the dynamics is periodic and described by a fictitious pendulum whose conserved energy coincides with $U_{\mathrm{osc}}$ Johns:2019izj ; Padilla-Gay:2021haz ; Fiorillo:2023mze ; Fiorillo:2023hlk . However, Eq. (6b) immediately reveals that the conservation breaks down for inhomogeneous settings
@@
For a numerical solution that extends to the nonlinear regime, we use periodic boundary conditions on a box of length $L=100$, divided in $N$ spatial bins. As initial conditions, we use $P^{z}_{\pm}(0)=\zeta_{\pm}=\pm 1/2$. These parameters imply that the range of unstable modes is delimited by $k_{1}=-2$ and $k_{1}=0$, the maximum growth rate obtains for $\overline{k}=-1$ and is $\overline{\gamma}=1$, and the initial oscillation energy is $U_{\mathrm{osc}}(0)=\zeta_{-}\zeta_{+}L=-25$. The transverse components $P^{x,y}_{\pm}$ are seeded with randomly sampled functions with a spatial dependence
@@
Figure 1 shows the solution for a single realization of initial seeds. The upper panel shows contours of the $\nu_{x}$-content of the beam initially occupied with $\nu_{e}$, as a function of $z$ and $t$. The lower panel shows $U_{\mathrm{osc}}(t)$ in units of $U_{\mathrm{osc}}(0)$. On the shown timescale, the range of unstable modes $-2<k<0$ grows nonlinear and then keeps oscillating, with beats between different modes causing an irregular pattern that in detail depends on the choice of seeds. In the longer run, these nonlinear modes feed higher-$k$ modes and flavor variations will obtain on ever smaller scales. Either way, the nonconservation of $U_{\mathrm{osc}}$ is a large effect, not a small correction.
@@
[…]
@@
For the interaction part, the calculations are much lengthier, so we here only outline the general strategy and report details in the Supplemental Material (SM). The commutator $[\mathcal{D},\mathcal{U}]$ produces strings of four creation and annihilation operators, or four-point correlation functions. The average of these combinations are evaluated again using the mean-field approximation. In the homogeneous limit, when taking the expectation value of $\langle a^{\dagger}_{1}a_{2}\rangle$, the momenta of state $1$ and $2$ must be equal. We go one step further and include the inhomogeneity to first order in ${\bf k}$. The final result, fully derived in the SM, is
@@
where $p=|{\bf p}|$ and $p^{\prime}=|{\bf p}^{\prime}|$. The correction to the quasi-particle energy can be obtained directly from the interaction Hamiltonian, taking the mean-field Hamiltonian expanded to first order in the gradients.
The new terms are intuitive if we interpret ${\sf\Omega}_{{\bf p},{\bf r}}$ as the renormalization of the quasiparticle energy induced by the interaction, which is in general a matrix in flavor space, analogous to the renormalization of quasiparticle energy in Landau’s theory of a Fermi liquid Landau:1958joj . This induces a renormalization in the group velocity ${\bm{\nabla}}_{\bf p}{\sf\Omega}^{(0)}_{{\bf p},{\bf r}}$ (only the zero order terms must be kept which already contain one gradient) and a weak force field $-{\bm{\nabla}}_{\bf r}{\sf\Omega}^{(0)}_{{\bf p},{\bf r}}$.
@@
is the number flux of neutrinos passing through a surface element of coordinate space at phase-space location $\{{\bf p},{\bf r}\}$, whereas ${\bf F}_{{\bf p},{\bf r}}$ is the same in momentum space, given by the same expression with ${\bm{\nabla}}_{\bf p}\to{\bm{\nabla}}_{\bf r}$. Crucially, this term couples the trace of the density matrix with the polarization vector; taking the trace, we find
@@
Thus, spatial gradients of the polarization vectors can feed $\nu$ kinetic energy. The order of magnitude of this term implies a net rate of energy gain or loss of order $d\epsilon_{\bf p}/dt\simeq\mu/\ell$. Over a timescale $\ell$, the energy that a $\nu$ can accumulate is of order $\mu\ll\epsilon_{\bf p}$, so we recover that this is a small effect relative to $\epsilon_{\bf p}$. It provides the missing channel by which kinetic and refractive energy can be traded. Even a small amount of energy $\mu$ lost (gained) from the large $\nu$ kinetic energy explains the change in the refractive energy, since $U_{\mathrm{osc}}/K\simeq\mu/\epsilon_{\bf p}$.
With the new kinetic equations, we can directly prove that to first order in $r_{\mu}$ and $r_{\ell}$, the total energy is indeed conserved because the kinetic energy $K$ is no longer conserved. After an integration by parts, we find
@@
Discussion.—We have shown that the usual kinetic equation for mixed neutrinos Eq. (1) is non-conservative in the inhomogeneous case. We have illustrated this point with a simple two-beam example, where the $\nu$$\nu$ interaction energy strongly changes, and we have also shown this effect analytically in the linear regime. Deriving small gradient terms beyond Eq. (1), we have shown that the refractive energy gained or lost is precisely traded with neutrino kinetic energy that usually is not followed. In our two-beam example, the monochromatic initial energy distribution develops small space-time dependent shifts that account for the missing energy. We have not tried to study this effect numerically because it requires many ${\bf p}$-modes around the original one, but we have proven energy conservation analytically.
The usual EOMs correctly account for flavor evolution, meaning the trace-free part of the density matrices often described by polarization vectors. On the other hand, the particle number in a given ${\bf p}$-mode is conserved, corresponding to the trace part of the EOM, where the left-hand side of Eq. (1) is a continuity equation. Therefore, the trace part of the EOM must be expanded to higher order in the gradients to capture nontrivial evolution. The reshuffling of neutrinos among ${\bf p}$-modes is a small effect relative to the large kinetic energies, yet precisely absorbs the missing refractive energy. Unless one wants to follow small modifications of the neutrino energy distribution, Eq. (1) remains the one to be studied.
We note that Eq. (1) conserves entropy and including the new gradient terms, this is also the case, i.e., entropy is conserved order by order in the gradient expansion (see Supplemental Material). On the practical level, the conservation of entropy, but not of energy, could be used to test numerical stability.
For the moment, our finding is primarily of conceptual interest: at least superficially, the roles of energy and entropy are somewhat exchanged relative to a naive expectation. This finding may shed new light on a recent proposal that the final outcome of fast conversions may be some sort of thermalized state Johns:2023jjt . The quasi-steady state generically observed in numerical simulations of fast conversions Padilla-Gay:2020uxa ; Zaizen:2023ihz ; Xiong:2023vcm ; Abbar:2023ltx ; Cornelius:2023eop ; Urquilla:2024bvf is only determined by the oscillatory part of the density matrix, yet its dynamics does not admit a separately conserved energy. Thus, it cannot separately thermalize, since it exchanges energy with the kinetic energy of neutrinos.
@@
We can now take the inverse Fourier transform by multiplying by $e^{-i{\bf k}\cdot{\bf r}_{1}}$ and summing over all ${\bf k}$. We perform the full manipulations only on the first of the four terms; the other ones are treated analogously. Introducing the notation ${\bf a}={\bf k}-{\bf q}$, $\mathbf{b}={\bf k}+{\bf q}$, we can write this term as
$\sum_{\frac{{\bf a}}{2},\frac{\mathbf{b}}{2},{\bf p}^{\prime},\gamma}e^{i\frac{\mathbf{b}}{2}\cdot({\bf r}_{2}-{\bf r}_{1})+i\frac{{\bf a}}{2}\cdot({\bf r}_{3}-{\bf r}_{1})}\rho_{\gamma\gamma}({\bf p}^{\prime},{\bf r}_{3})\rho_{\alpha\beta}\left({\bf p}-\frac{{\bf a}}{4},{\bf r}_{2}\right)\left({\bf p}+\frac{{\bf q}}{2},{\bf p}+\frac{{\bf k}}{2}\right)\left({\bf p}^{\prime}+\frac{{\bf a}}{4},{\bf p}^{\prime}-\frac{{\bf a}}{4}\right).$
In this expression, we must now perform the expansion for $|{\bf k},{\bf q}|\ll|{\bf p},{\bf p}^{\prime}|$. For the spinor part, this requires using the explicit form of the spinors expanded to first order in the momenta ${\bf a}$ and $\mathbf{b}$. The result depends on the mutual orientation of the vectors ${\bf p}$ and ${\bf p}^{\prime}$. Thus, we introduce a set of three orthonormal directions

### v2

@@ Abstract
Flavor-dependent neutrino transport is described by a well-known kinetic equation for occupation-number matrices in flavor space. However, in the context of fast flavor conversion, we identify an unforeseen predicament: the pivotal self-induced exponential growth of small inhomogeneities strongly violates conservation of neutrino-neutrino refractive energy. We prove that it is traded with the huge reservoir of neutrino kinetic energy through gradients of neutrino flavor coherence (the off-diagonal piece of the flavor density matrix) and derive the missing gradient terms. The usual equations remain sufficient to describe flavor evolution, at the cost of renouncing energy conservation, which cannot play any role in explaining the numerically observed final state.
Introduction.—Flavor evolution in a dense neutrino environment is often described by a kinetic equation for the flavor density matrices $\rho({\bf p},{\bf x},t)$, where the diagonal entries are the occupation numbers Dolgov:1980cq ; Rudsky ; Sigl:1993ctk ; Sirera:1998ia ; Yamada:2000za ; Vlasenko:2013fja ; Cirigliano:2014aoa ; Volpe:2013uxl ; Serreau:2014cfa ; Kartavtsev:2015eva ; Fiorillo:2024wej . In the homogeneous case, they are defined as the expectation values of the number-operator matrices ${\cal D}_{\alpha\beta}=a^{\dagger}_{\beta,{\bf p}}a_{\alpha,{\bf p}}$ with the flavor indices $\alpha$ and $\beta$ and $a^{\dagger}_{\alpha,{\bf p}}$ the creation operator for a left-handed neutrino of flavor $\alpha$ with momentum ${\bf p}$. In the limit of vanishing neutrino masses and flavor mixing, nontrivial flavor evolution can still arise through classical instabilities that would engender strong flavor correlations Sawyer:2015dsa ; Chakraborty:2016lct ; Izaguirre:2016gsx ; Tamborra:2020cul ; Richers:2022zug . Dubbed fast flavor evolution, this subject has been intensely studied in view of practical consequences for neutrino flavor transport in stellar core collapse and binary neutron star mergers Xiong:2022vsy ; Fernandez:2022yyv ; Ehring:2023abs ; Ehring:2023lcd ; Nagakura:2023mhr ; Nagakura:2023xhc ; Zaizen:2023wht ; Grohs:2023pgq ; Abbar:2023zkm ; Akaho:2023brj ; Froustey:2023skf ; Nagakura:2023wbf .
Predicting the consequences of fast conversions is thus a central question. In tackling this problem, a key role must be played by conserved quantities, which are the only exact guide into the final state reached after conversions. One such quantity must be energy. However,we show that the usual equations of motion (EOMs) do not conserve energy. In their often-used form, they are
@@
where ${\sf H}_{{\bf p},{\bf r},t}=\sqrt{2}G_{\rm F}\sum_{{\bf p}^{\prime}}\rho_{{\bf p}^{\prime},{\bf r},t}\,(1-\cos\theta_{{\bf p},{\bf p}^{\prime}})$ is a matrix driving flavor evolution by neutrino-neutrino refraction. The advection term proportional to ${\bf v}={\bf p}/|{\bf p}|$ quantifies a drift in coordinate space caused by inhomogeneities. This term allows for ${\bf p}$-dependent instabilities, so even an initially homogeneous system can develop flavor disturbances growing exponentially and drifting to ever smaller scales Duan:2014gfa ; Mirizzi:2015fva ; Bhattacharyya:2020jpj ; Johns:2020qsk ; Tamborra:2020cul ; Richers:2021nbx ; Bhattacharyya:2022eed .
@@
Interaction energy.—We first define the $\nu$$\nu$ interaction energy. The starting point is the many-body Hamiltonian $\mathcal{H}=\mathcal{H}_{0}+\mathcal{U}$, where $\mathcal{H}_{0}=\sum_{\alpha,{\bf p}}\epsilon_{\bf p}a^{\dagger}_{\alpha,{\bf p}}a_{\alpha,{\bf p}}$ is the kinetic energy operator with $\epsilon_{\bf p}=|{\bf p}|$ in the massless limit. Moreover, the interaction energy is
@@
The energy of the system is the expectation value of $\mathcal{H}$ over the quantum state. In particular, the kinetic energy is $K=\langle\mathcal{H}_{0}\rangle=\int d^{3}{\bf r}\sum_{\bf p}\epsilon_{\bf p}\mathrm{Tr}(\rho_{\bf p})$. For the interaction energy, we need to determine the average value of a string of four creation and annihilation operators. We use mean field approximation, where the quantum state is assumed to be the product of single-particle states. The expectation value factorizes as $\langle a^{\dagger}_{1}a_{2}a^{\dagger}_{3}a_{4}\rangle=\langle a^{\dagger}_{1}a_{2}\rangle\langle a^{\dagger}_{3}a_{4}\rangle+(\delta_{2,3}-\langle a^{\dagger}_{3}a_{2}\rangle)\langle a^{\dagger}_{1}a_{4}\rangle$. In this way we define a potential energy of interaction $U=\langle\mathcal{U}\rangle$. For a homogeneous system, it is
@@
where we neglect an irrelevant (infinite) renormalization of the neutrino chemical potential.
@@
Different from a neutrino gas subject to collisions with matter (see, e.g., Refs. Johns:2021qby ; Nagakura:2022qko ; Xiong:2022zqz ; Liu:2023pjw ; Lin:2022dek ; Johns:2022yqy ; Shalgar:2022rjj ; Padilla-Gay:2022wck ; Fiorillo:2023ajs ; Johns:2023xae ), where energy is not conserved anyway, Eqs. (6) describe a closed system that admits a conserved energy. The kinetic part $K=\int d^{3}{\bf r}\sum_{\bf p}\epsilon_{\bf p}P^{0}_{\bf p}$ is seen to be conserved from Eq. (6a). Also the nonoscillatory part of the interaction energy is separately conserved. $U_{\mathrm{osc}}=(\sqrt{2}G_{\rm F}/4)\int d^{3}{\bf r}\,({\vec{P}}_{0}^{2}-{\vec{P}}_{1}^{2})$, which follows from Eq. (5), must then be conserved as well. In the homogeneous case, this is indeed the case; the dynamics is periodic and described by a fictitious pendulum whose conserved energy coincides with $U_{\mathrm{osc}}$ Johns:2019izj ; Padilla-Gay:2021haz ; Fiorillo:2023mze ; Fiorillo:2023hlk . However, Eq. (6b) immediately reveals that the conservation breaks down for inhomogeneous settings
@@
For a numerical solution that extends to the nonlinear regime, we use periodic boundary conditions on a box of length $L=100$, divided in $N$ spatial bins. As initial conditions, we use $P^{z}_{\pm}(0)=\zeta_{\pm}=\pm 1/2$. These parameters imply that the range of unstable modes is delimited by $k_{1}=-2$ and $k_{2}=0$, the maximum growth rate obtains for $\overline{k}=-1$ and is $\overline{\gamma}=1$, and the initial oscillation energy is $U_{\mathrm{osc}}(0)=\zeta_{-}\zeta_{+}L=-25$. The transverse components $P^{x,y}_{\pm}$ are seeded with randomly sampled functions with a spatial dependence
@@
Figure 1 shows the solution for a single realization of initial seeds. The upper panel shows contours of the $\nu_{x}$-content of the beam initially occupied with $\nu_{e}$, as a function of $z$ and $t$. The lower panel shows $U_{\mathrm{osc}}(t)$ in units of $U_{\mathrm{osc}}(0)$. On the shown timescale, the range of unstable modes $-2<k<0$ grows nonlinear and then keeps oscillating, with beats between different modes causing an irregular pattern that in detail depends on the choice of seeds. In the longer run, these nonlinear modes feed higher-$k$ modes and flavor variations will obtain on ever smaller scales. If we had used a larger box, or equivalently averaged over more than one realization of the initial conditions, the amplitude of the oscillations in the final state would shrink. Nevertheless, the bulk of the nonconservation of $U_{\mathrm{osc}}$ happens in the initial phase, and is clearly a large effect.
@@
Inhomogeneous kinetic equations.—The lack of refractive-energy conservation questions the validity of the traditional EOMs. The strategy for their derivation is a perturbative expansion in three small parameters: (i) The mass-to-energy ratio $r_{m}=m_{\nu}/\epsilon_{\bf p}$, where $m_{\nu}$ is the neutrino mass, but in our fast flavor limit we do not worry about it. (ii) The fractional refractive energy shift $r_{\mu}=\mu/\epsilon_{\bf p}$, where $\mu=\sqrt{2}G_{\rm F}n_{\nu}$ is a scale for the $\nu$$\nu$ interaction energy. For $T\simeq 5$ MeV in the decoupling region of a supernova, $n_{\nu}\simeq 10^{33}$ cm-3, $\mu\simeq 0.2$ meV, and $r_{\mu}\simeq 10^{-11}$ and thus indeed very small. (iii) The scale of density variations $\ell$ in units of the neutrino de Broglie wavelength $r_{\ell}=(\ell|{\bf p}|)^{-1}\ll 1$. Density variations are on scales much larger than $|{\bf p}|^{-1}$. The EOMs are typically derived to lowest order in all of these small ratios, notably neglecting terms of the order of $r_{\ell}r_{\mu}$, which causes the nonconservation of energy.
@@
For the interaction part, we here outline the general strategy and report details in the Supplemental Material (SM). The commutator $[\mathcal{D},\mathcal{U}]$ produces strings of four creation and annihilation operators, or four-point correlation functions. The average of these combinations are evaluated again using the mean-field approximation. In the homogeneous limit, when taking the expectation value of $\langle a^{\dagger}_{1}a_{2}\rangle$, the momenta of state $1$ and $2$ must be equal. We go one step further and include the inhomogeneity to first order in ${\bf k}$. The final result, fully derived in the SM, is
@@
where $p=|{\bf p}|$ and $p^{\prime}=|{\bf p}^{\prime}|$.
${\sf\Omega}_{{\bf p},{\bf r}}$ is the renormalized quasiparticle energy, in general a matrix in flavor space, analogous to the renormalized quasiparticle energy in Landau’s theory of a Fermi liquid Landau:1958joj . This induces a renormalization in the group velocity ${\bm{\nabla}}_{\bf p}{\sf\Omega}^{(0)}_{{\bf p},{\bf r}}$ (only the zero order terms must be kept which already contain one gradient) and a weak force field $-{\bm{\nabla}}_{\bf r}{\sf\Omega}^{(0)}_{{\bf p},{\bf r}}$.
@@
is the number flux of neutrinos passing through a surface element of coordinate space at phase-space location $\{{\bf p},{\bf r}\}$, as seen from the structure analogous to a continuity equation, whereas ${\bf F}_{{\bf p},{\bf r}}$ is the same in momentum space, given by the same expression with ${\bm{\nabla}}_{\bf p}\to{\bm{\nabla}}_{\bf r}$. This term couples the trace of the density matrix with the polarization vector; taking the trace, we find
@@
Thus, spatial gradients of the polarization vectors can feed $\nu$ kinetic energy. This term implies a net rate of energy gain or loss of order $d\epsilon_{\bf p}/dt\simeq\mu/\ell$. Over a timescale $\ell$, the energy that a $\nu$ can accumulate is of order $\mu\ll\epsilon_{\bf p}$, so we recover that this is a small effect relative to $\epsilon_{\bf p}$, but large relative to the interaction energy. It provides the missing channel by which kinetic and refractive energy can be traded. Even a small amount of energy $\mu$ lost (gained) from the large $\nu$ kinetic energy explains the large change in the refractive energy, since $U_{\mathrm{osc}}/K\simeq\mu/\epsilon_{\bf p}$.
With the new kinetic equations, we can directly prove that to first order in $r_{\mu}$ and $r_{\ell}$, the total energy is conserved. After integrating by parts, we find
@@
Discussion.—We have shown that the usual kinetic equation for mixed neutrinos Eq. (1) is non-conservative in the inhomogeneous case. We have illustrated this point with a simple two-beam example, where the $\nu$$\nu$ interaction energy strongly changes, and we have also shown this effect analytically in the linear regime. Deriving missing gradient terms beyond Eq. (1), we have shown that the refractive energy gained or lost is precisely traded with neutrino kinetic energy that usually is not followed. In our two-beam example, the monochromatic initial energy distribution develops small space-time dependent shifts that account for the missing energy. We have not tried to study this effect numerically because it requires many ${\bf p}$-modes around the original one, but we have proven energy conservation analytically.
The usual EOMs correctly account for flavor evolution, meaning the trace-free part of the density matrices often described by polarization vectors. On the other hand, the particle number in a given ${\bf p}$-mode is conserved, corresponding to the trace part of the EOM, where the left-hand side of Eq. (1) is a continuity equation. Therefore, the trace part of the EOM must be expanded to higher order in the gradients to capture nontrivial evolution. The reshuffling of neutrinos among ${\bf p}$-modes is a small effect relative to the large kinetic energies, yet precisely absorbs the missing refractive energy.
These novel terms come from the gradients of the neutrino self-energy. With hindsight, that they would affect neutrino evolution, is obvious from the viewpoint of physical kinetics, and implicitly present in previous formal derivations Sigl:1993ctk ; Rudsky ; Yamada:2000za ; Vlasenko:2013fja ; Cirigliano:2014aoa , but usually assumed to be a small effect. Our new insight is that fast conversions spontaneously break homogeneity, magnifying the gradients of the trace-free density matrix, making these terms large enough to explain completely the apparent non-conservation of energy.
We note that Eq. (1) conserves entropy and including the new gradient terms, this is also the case, i.e., entropy is conserved order by order in the gradient expansion (see Supplemental Material). On the practical level, the conservation of entropy, but not of energy, could be used to test numerical stability. Even more importantly, in making predictions on the final state induced by conversions, energy conservation cannot be used in practice, unless the new terms are kept.
Our finding may shed new light on a recent proposal that the final outcome of fast conversions may be some sort of thermalized state Johns:2023jjt . The quasi-steady state generically observed in numerical simulations of fast conversions Padilla-Gay:2020uxa ; Zaizen:2023ihz ; Xiong:2023vcm ; Abbar:2023ltx ; Cornelius:2023eop ; Urquilla:2024bvf is only determined by the oscillatory part of the density matrix, yet its dynamics does not admit a separately conserved energy. Thus, it cannot separately thermalize, since it exchanges energy with the kinetic energy of neutrinos.
@@
We can now take the inverse Fourier transform by multiplying by $e^{-i{\bf k}\cdot{\bf r}_{1}}$ and summing over all ${\bf k}$. We perform the full manipulations only on the first of the four terms; the other ones are treated analogously. Introducing the notation ${\bf a}={\bf k}-{\bf q}$, ${\bf b}={\bf k}+{\bf q}$, we can write this term as
$\sum_{\frac{{\bf a}}{2},\frac{{\bf b}}{2},{\bf p}^{\prime},\gamma}e^{i\frac{{\bf b}}{2}\cdot({\bf r}_{2}-{\bf r}_{1})+i\frac{{\bf a}}{2}\cdot({\bf r}_{3}-{\bf r}_{1})}\rho_{\gamma\gamma}({\bf p}^{\prime},{\bf r}_{3})\rho_{\alpha\beta}\left({\bf p}-\frac{{\bf a}}{4},{\bf r}_{2}\right)\left({\bf p}+\frac{{\bf q}}{2},{\bf p}+\frac{{\bf k}}{2}\right)\left({\bf p}^{\prime}+\frac{{\bf a}}{4},{\bf p}^{\prime}-\frac{{\bf a}}{4}\right).$
In this expression, we must now perform the expansion for $|{\bf k},{\bf q}|\ll|{\bf p},{\bf p}^{\prime}|$. For the spinor part, this requires using the explicit form of the spinors expanded to first order in the momenta ${\bf a}$ and ${\bf b}$. The result depends on the mutual orientation of the vectors ${\bf p}$ and ${\bf p}^{\prime}$. Thus, we introduce a set of three orthonormal directions